OpenStudy (anonymous):

The volume of a sphere can be given by V = 4.18879r3. You have to design a spherical container that will hold a volume of 55 cubic inches. What should the radius of the container be?

OpenStudy (unklerhaukus):

\[V_{\text{ sphere}}=\frac 43\pi r^3\approx4.18879r^3\] \[r=\sqrt[3]{\frac 3{4\pi}V}\approx \sqrt[3]\frac V{4.18879}\] \[V=55 \left[\text{ in}^3\right]\] \[r=\sqrt[3]{\frac 3{4\pi}55 \left[\text{ in}^3\right]}\approx \sqrt[3]\frac {55 \left[\text{ in}^3\right]}{4.18879}=\sqrt[3]{\frac{55}{4.18879}}[\text {in}]\]

OpenStudy (anonymous):

Ohemgee thank you bunches :) Medal for you!!!! :D